OpenStudy (anonymous):

A 10 foot ladder is leaning against a house. The bottom of the ladder is 5 feet from the building. How high is the top of the ladder? Round to the nearest hundredth

OpenStudy (saifoo.khan):



OpenStudy (saifoo.khan):

using pythagoras.

OpenStudy (saifoo.khan):

\[10^2 - 5^2 = h^2\]h = 8.66